A 0.3 kg horizontal mass-spring system undergoes SHM. The period is 0.60 s. When the mass passes its equilibrium position it has a speed of 1.5 m/s. Assuming the surface is frictionless, what is the total mechanical energy and the amplitude of the system?
Answer #1

mechanical energy of system

U = 0.5 m v^2 = 0.5* 0.3* 1.5* 1.5 = 0.03375 J

=====

also

v = A (2 pi / T)

1.5 = A ( 2 pi / 0.6)

A = 0.1433 m

=====

Comment before rate in case any doubt, will reply for sure.. goodluck